A curve has the equation y=3x^2-5x+4. Find the gradient of the curve at the point (2,6)


1
Expert's answer
2021-05-24T16:12:10-0400

y=3x^2-5x+4 ...................Equation(1)


Gradient of a curve at any point (x, y) is given by


m = "\dfrac{dy}{dx}"


Differentiating Equation(1) with respect to x we have,



"\dfrac{dy}{dx}" = 6x - 5


"\lparen\dfrac{dy}{dx}\rparen" at x = 2 is equal to


m = 6"*" 2 - 5

m = 7